OpenStudy (anonymous):

can someone help in solving for t? t^6+4t^3=45 ......i got 5 and -9 as my answer but when i put it as my answer its saying its wrong :(

OpenStudy (anonymous):

you should have \[t^3=9\] or \[t^3=-3\]

OpenStudy (anonymous):

can you show me the steps you took to get to this please?

OpenStudy (anonymous):

wait its saying those answers are wrong too...

OpenStudy (anonymous):

ok maybe that was wrong. lets start at the beginning. first write \[t^6+4t^3-45=0\] then factor as \[(t^3+9)(t^3-5)=0\] so \[t^3=-9\] or \[t^3=5\] and therefore \[t=\sqrt[3]{-9}\] or \[t=\sqrt[3]{5}\]

OpenStudy (anonymous):

thank you both!